solve the system by graphing
y=-9-2x
x+2y=-12

Answer by MisfortunateMonstrosity(1)

First of all, you need to make both equations slope-intercept equations.
Y=Mx+B
y is the y, m is the slope, x is the x, and b is the y intercept (the number on the y line)
What is equal to Y=-9-2X that would match the Slope Intercept form?
Y=-2X-9.
Same question for next equation.
X+2Y=-12
2Y=-X-12.
Y should be Y by itself, so:
Y=-.5X-6 (You could just do 1/2 instead of .5)
Now you graph both of the equations
Y=-.5X-6 and Y=-2X-9
Use the Y intercepts; -6 and -9, put them both on the Y line in the graph.
Then you use the slope to make the lines in the graph.
For the first equation (Y=-.5-6), you have to graph 0,-6, and then you use the slope to make it into a line. Remember that -1/2 is rise/run, so you go up 1 time, and then since it is a negative you have to go left instead of right, so you go left twice.
Do the same for the second equation:
Graph (0,-9), and use the slope -2/1 to graph a line.
Use a ruler for the lines, and keep drawing both lines until they touch. Where ever they touch is your answer.
Try it, and if you did it correctly, (-2,-5) should be your answer.